OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line containing the median of the trapezoid whose vertices are R(-1, 5) , S(1, 8), T(7, -2), and U(2, 0).

OpenStudy (anonymous):

@jinxhead20 Can you help? Please?

OpenStudy (anonymous):

y - y1 = m(x - x1), with m = -5/3 and (x1, y1) = (0, 6.5) y - 6.5 = (-5/3)(x - 0) y - 6.5 = -5x/3 5x/3 + y = 6.5 (5/3)x + y = 6.5 or if you prefer multiply both sides by 6 to get rid of the denominator and the decimal 10x + 6y = 39

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

Yes! thank you :)

OpenStudy (anonymous):

can you press the Best Response button right next to my answer please?

OpenStudy (anonymous):

done:)

OpenStudy (anonymous):

thanks